OpenStudy (anonymous):

an appliance store decreases the price of a 19 inch television set 20% to a sale price of 507.20. what was the original price? thanks <3

OpenStudy (anonymous):

original price=507.20+507.20*.02=517.34

OpenStudy (anonymous):

\[\frac{507.20}{0.8}=634.00 \]

OpenStudy (anonymous):

thankss:) where did you get the 0.8 from?:)

OpenStudy (anonymous):

100% -20% = 80% or 0.8

OpenStudy (anonymous):

jesus crist ! thanks so much:p

OpenStudy (anonymous):

Your welcome.